OpenStudy (anonymous):

sin7.5= cos7.5= tan7.5= cot7.5=

OpenStudy (callisto):

7.5 in rad. or in degree?

OpenStudy (anonymous):

no with solutions

OpenStudy (callisto):

i meant the question

OpenStudy (anonymous):

Do you want the answers in radians or in degrees?

OpenStudy (anonymous):

for example: sin15=2cos7.5\[\times\]sin7.5

OpenStudy (callisto):

sin15=/= 2cos7.5!!!

OpenStudy (anonymous):

no for example it means you solve it

OpenStudy (anonymous):

do you understand?

OpenStudy (callisto):

no!

OpenStudy (anonymous):

I do not want,I solve it and send for you

OpenStudy (anonymous):

may you solve it?

OpenStudy (callisto):

i thought you were saying that sin15 = 2sin7.5cos 7.5 , right?

OpenStudy (anonymous):

No No i said for example

OpenStudy (anonymous):

It means This way I can solve

OpenStudy (callisto):

you're not solving.. you're just expressing it in a more complicated way..

OpenStudy (anonymous):

Yes Hey freind I'm From Iran and PM is here and I go eat lunch

OpenStudy (anonymous):

see you later for 20min or 1 day may solve it? thank you very much

OpenStudy (callisto):

sin7.5= 2sin3.75cos3.75 cos7.5=2(cos3.75)^2 - 1 tan7.5= (2tan3.75)/[1-(tan3.75)^2] cot7.5= 1/tan7.5 =[1-(tan3.75)^2]/ (2tan3.75) If that's what you want

OpenStudy (anonymous):

yes yes thank you very much see you lator and I go eat lunch because I'm very hungry